Golden Rectangle. Greeks to get shapes that are pleasing to the eye in art and architecture was the Golden Rectangle. If a square is removed from one end of a Golden Rectangle, the sides of the remaining rectangle are proportional to the original rectangle. So the length and width of the original rectangle satisfy L over W = W over L minus W. If the length of a Golden Rectangle is 10 meters, then what is its width?
Golden painting. An artist wants her painting to be in the shape of a golden rectangle. If the length of the painting is 36 inches, then what should be the width?

Answer by richwmiller(17219)

L/W=W/(L-W)
10/W=W/(10-W)
36/W=W/(36-W)
or is it
L/W=(W/L)-W
10/W=(W/10)-W
36/W=(W/36)-W
These return complex numbers for W.